David has 50 employees. Blue collar workers make $600 per week and white collar workers make $1000 per week. David's payroll is $42,000. David has how many white collar workers?
Answer by NancyLam(35)

I don't do add or subtrac both side it take to long and make confusing too, I do only one side it quick and easy, you can do both side if you want
Let x= blue , y= White
$600x+$1000y=$42000
x+y=50
y= 50-x
600x+1000(50-x)=42000 (subtitution y=50-x)
600x+50000-1000x =42000 (mutipy)
-400x =42000-50000 (simplify)
x =-8000/-400
x = 20
x+y+=50
20+y=50
y = 50-20
y = 30
Check: y=30, x=20
$600(20)+1000(30)=$12000+$30000=$42000